An electrical firm manufactures light bulbs that have a length of life that is approximately normally

distributed with a standard deviation of 40 hours. If a sample of 30 bulbs has an average life of 780

hours, find a 96% confidence interval for the population mean of all bulbs produced by this firm


1
Expert's answer
2021-10-18T15:14:14-0400

Solution:

The critical value for z distribution for lower tail and at 96% level of significance is 2.05.

The 96% confidence interval for the population mean of all bulbs produced by this firm is obtained below:

"\bar x\pm z_{\alpha/2}(\dfrac{\sigma}{\sqrt n})
\\=780\pm2.05(\dfrac{40}{\sqrt{30}})
\\=780\pm14.97
\\=(765.03,794.97)"